OpenStudy (anonymous):

is squarerootof(x-5)/2 a function??

OpenStudy (anonymous):

you mean, \[\sqrt{f(x-5)/2}\] ? If so, yes.

OpenStudy (anonymous):

you mean \[f(x)=\sqrt{\frac{x-5}{2}}\]? it is a function for sure